OpenStudy (anonymous):

x^4 + 1 ------- = ? x^2 - 1

OpenStudy (ash2326):

Are you sure the question isn't? \[\frac{x^4-1}{x^2-1}\]

OpenStudy (anonymous):

yes, thats it

OpenStudy (ash2326):

We know \[a^2-b^2=(a+b)(a-b)...............................(1)\] here we have, \[\frac{x^4-1}{x^2-1}\] Now we'll use (1) to factor the numerator \[x^4-1=(x^2-1)(x^2+1)\] so we have \[\frac{(x^2-1)(x^2+1)}{x^2-1}\] Let's cancel the common terms from numerator and denominator \[\frac{\cancel {(x^2-1)}(x^2+1)}{\cancel{x^2-1}}\] We get finally \[{x^2+1}\]

OpenStudy (anonymous):

what if we had an \[x^{4}+1\] in the numerator? could you solve that for me?

OpenStudy (ash2326):

Then it can't be reduced further:(

OpenStudy (anonymous):

alright then... thanks for the help! :)

OpenStudy (ash2326):

You're welcome:) Welcome to Open Study:)